OpenStudy (anonymous):

Which value of x makes the following equation true? 3(x - 5) - 16 = 2(x - 8) + 3 18 26 13 10

OpenStudy (campbell_st):

might be easiest to just solve it add 16 to both sides 3(x - 5) = 2(x - 8) + 19 now distribute \[3x - 15 = 2x -16 + 19\] so its 3x - 15 = 2x + 3 hope that's a start that makes sense

OpenStudy (anonymous):

im still confused